OpenStudy (anonymous):

which is longer' 5 miles or 10 kilometers

OpenStudy (texaschic101):

1 km = 0.6214 miles 1/0.6214 = 10/x -- (1 km to 0.6214 m = 10 km to x miles) cross multiply (1)(x) = (0.6214)(10) x = 6.214 so this tells me that 10 km = 6.214 miles so what do you think the answer is ?